Ryan defeated Tom in an election for reeve of the township. A total of 6788 votes were cast. If 100 voters had switched their votes from Ryan to Tom, then Tom would have won by 64 votes. How many votes did each candidate receive?
Answer by ikleyn(52786)

.

    R + T = 6788

    R - 100 = T + 100  - 64.


Simplify


    R + T = 6788

    R - T =  136.


Add the last two equations. You will get


    2R = 6788 + 136 = 6924  ====>  R = 6924/2 = 3462.


ANSWER.  Ryan received  3462 votes.  Tom received the rest,  6788 - 3462 = 3326 votes.

Solved.